OpenStudy (anonymous):

Can somebody prove the Dot Product Theorem for me? Can somebody also explain why are a and b are in absolute value signs and provide a visual representation of the theorem if possible?

OpenStudy (anonymous):

\[ab = \left| a \right| \left| b \right| \cos \theta \]

OpenStudy (anonymous):

\[a dotb=a_{x} b_{x}+a_{y} b_{y}+a_{z} b_{z}\] This is what Gibbs defined it to be a long time ago. b(x)=abs(a) cosa(lpha) b(y)=asina a(x)=bcosb a(y)=bsinb cos(b-a)=cos(theta)=cosbcosa-sinbsina (see euler equation) \[a_{x} b_{x}+a_{y} b_{y}=acos(a)bcos(b)+asin(a)bsin(b)=abs(a)abs(b)cos(b-a)\]Sorry for the confusing notationing.